Early in its fiscal year ending December 31, 2018, San Antonio Outfitters finalized plans to expand operations. The first staInitial Value Assets Land Land improvements Building Equipment Furniture & fixtures

Interest expense
Answer #1
Please hit LIKE button if this helped. For any further explanation, please put your query in comment, will get back to you.
Requirement 1
Land
Purchase price (determined below) $                                                                   1,055,300
Closing costs $                                                                         37,000
Removal of old building $                                                                         87,000
Clearing and grading $                                                                         67,000
$                                                                   1,246,300
Purchase price of land:
      Cash paid $                                                                       370,000
      Value of note (From below working) $                                                                       685,300
$                                                                   1,055,300
Present value of note payment:
PV = $770,000 (.89000) $                                                                       685,300
Present value of $1: n = 2, i = 6%
Land improvements
Parking lot and landscaping $                                                                       370,000
Building
Construction expenditures:  
5/1 $                                                                   3,750,000
7/1 $                                                                   2,350,000
9/1 $                                                                   1,920,000
10/1 $                                                                   2,820,000
    Total expenditures $                                                                 10,840,000
Interest capitalized (determined below) $                                                                       301,750
      Total cost of building $                                                                 11,141,750
Average accumulated expenditures:
May 31, 2018 $3,750,000 x   6/6 = $                  3,750,000
July 30, 2018 2,350,000 x   3/6 = $                  1,175,000
September 1, 2018 1,920,000 x   2/6 = $                      640,000
October 1, 2018 2,820,000 x   1/6 = $                      470,000
$                  6,035,000
Interest capitalized:
$6,035,000 x6% x 6/12 = $                                                                       301,750
Equipment and furniture and fixtures
      Initial
Percent of Total     Valuation
Fair Value Fair Value % x $770,000
Equipment $                                                                       522,000                                 0.60 $            462,000
Furniture & fixtures $                                                                       348,000                                 0.40 $            308,000
Totals $                                                                       870,000                                 1.00 $            770,000
Initial valuation:
Equipment $                                                                       462,000
Furniture & fixtures $                                                                       308,000
Requirement 2
Interest expense:
Note issued to purchase land and building,
     $685,300 x 6% x 9/12 = $                                                                         51,398
Construction loan, $6,100,000 x 10% x 8/12 $                                                                       244,000
Long-term note, $3,700,000 x 7% $                                                                       259,000
Long-term bonds, $5,700,000 x 4% $                                                                       228,000
      Total $                                                                       782,398
Less: Interest capitalized (determined above) $                                                                     -301,750
Interest expense $                                                                       480,648